# Towards LifeSpan Cognitive Systems

### v1

@@ 7 Our Suggested Paradigm of LSCS
Existing technologies individually address some aspects of the challenges associated with Abstracting and Merging Experiences and Long-term Retention and Accurate Recall, but none can handle both sets of issues comprehensively. To address these limitations, we propose a new paradigm of LifeSpan Cognitive System (LSCS) that integrates and combines these methods to create a more robust solution. As depicted in Figure 2, the LSCS operates in two key phases: absorbing experiences and generating responses to environmental queries, with the detailed descriptions shown below:
@@ 7.1 Absorbing Experiences
Non-semantic Information Storage: Certain types of data, such as phone numbers or some hard-to-remember names, are challenging for typical memory models to store effectively. These pieces of non-semantic information, which lack comprehensive contextual meaning but are critical for precision, are stored in a dedicated knowledge base. The knowledge base acts as a repository for factual, easily forgotten information that LSCS can access as needed.
Semantic Information Abstraction: Experiences are also abstracted to capture their core meaning or “essence,” such as a general impression of a person, an emotional response to an event, or a rough memorization of some events that happened before. These abstracted memories, which may lack fine details, are stored in the external memory module. This module allows for memory consolidation and combination, leading to the creation of new insights and deeper understandings through a process akin to memory integration and fusion.
@@
As experiences are abstracted through these levels, the degree of compression increases, transitioning from raw text to structured knowledge, and eventually to internalized model parameters.
@@ 7.2 Generating Responses to Environmental Queries
Response Generation: With all necessary inputs in place, the large language model (potentially a multi-modal one as the experiences can include images, speech, videos as well [118]). takes the context as the input and generate a response according to the context and the model parameters. This step includes reading the extracted information from the knowledge base, the memory module, and the model parameters to provide a complete, context-aware response to the environmental query.
7.3 Summary of the Proposed Paradigm
In the above-mentioned processes, both the knowledge base and memory module are utilized, allowing us to adopt the technologies described in Saving $\mathcal{E}$ into Knowledge Bases for Retrieval and Saving $\mathcal{E}$ into Explicit Memory. Since the system incorporates a large language model, the techniques outlined in Saving $\mathcal{E}$ into Model Parameters are also applicable. During response generation, the model must handle extensive contextual inputs, which can be long enough to require the large language model to have long-context processing capabilities. By leveraging these technologies, we can enable Abstraction and Experience Merging, where the knowledge base, external memory module, and model parameters can all store abstracted versions of experiences. Then with the presence of both the knowledge base and external memory module, the system is capable of achieving Long-term Retention and Accurate Recall via querying these modules.
@@ 8 Conclusion
In this paper, we propose the concept of a LifeSpan Cognitive System (LSCS), which aims to manage rapid, incremental learning while retaining the ability to recall previous experiences. To achieve LSCS, we argue there are two major challenges: Abstraction and Experiences Merging, Long-term Retention and Accurate Recalling. Existing technologies have the potential to partially solve these challenges. Based on our defined metric, Storage Complexity, for saving past experiences ($\mathcal{E}$), we categorize existing technologies into four distinct classes. We summarize various technologies within each category the strengths and weaknesses of the methods in each category. We argue that achieving an LSCS requires a nuanced approach that integrates multiple strategies to address the two challenges. To this end, we propose a new paradigm for LSCS. We hope this paper provides valuable insights into LSCS and encourages further efforts to build more human-like cognitive systems.

### v2

@@ 7 Our Proposed Instantiation of LSCS
Existing technologies individually address some aspects of the challenges associated with Abstracting and Merging Experiences and Long-Term Retention and Accurate Recall, but none can handle both sets of issues comprehensively. To address these limitations, we propose a new instantiation that integrates and combines these methods to create a more robust solution. As depicted in Figure 2, our instantiation operates in two key phases: absorbing experiences and generating responses to environmental queries, with the detailed descriptions shown below:
@@ 7.1 Absorbing Experiences
Non-semantic Information Storage: Certain types of data, such as phone numbers or some hard-to-remember names, are challenging for typical memory models to store effectively. These pieces of non-semantic information, which lack comprehensive contextual meaning but are critical for precision, are stored in a dedicated knowledge base. The knowledge base acts as a repository for factual, easily forgotten information that LSCS can access as needed. Just as humans may use notebooks to store specific pieces of information they cannot reliably and easily retain in their memory (e.g., phone numbers, exact names, addresses), an LSCS could have an external, easily accessible storage component serving a similar role. For instance, frequently updated, high-fidelity information like contact details, passcodes, or historical records (generally “verbatim” facts that are cumbersome to store in model parameters) would be stored in such a dedicated external knowledge base, thus we add the term “Notepad” in Figure 2(a) to highlight the role of the knowledge base.
Semantic Information Abstraction: Experiences are also abstracted to capture their core meaning or “essence,” such as a general impression of a person, an emotional response to an event, or a rough memorization of some events that happened before. These abstracted memories, which may lack fine details, are stored in the external memory module. This module allows for memory consolidation and combination, leading to the creation of new insights and deeper understandings through a process akin to memory integration and fusion. The distinction between semantic information and non-semantic information is that non-semantic information is generally much harder to remember for humans. Humans can memorize what happened in the movie (which falls under the semantic information category) after watching it once and can recall the plot after a long time but may struggle to remember some detailed information such as phone numbers, specific names that do not have any semantic meaning (we call them non-semantic information). Thus for our instantiation, we also handle these two types of information differently.
@@
As experiences are abstracted through these levels, the degree of compression increases, transitioning from raw text to structured knowledge, and eventually to internalized model parameters. During this process, conflicts between new and existing knowledge may inevitably arise. Ideally, a memory model designed to resemble human memory could effectively merge similar experiences over time, while a dynamically updated knowledge graph could help resolve conflicts and integrate new information seamlessly.
@@ 7.2 Generating Responses to Environmental Queries
Response Generation: With all necessary inputs in place, the large language model (potentially a multi-modal one as the experiences can include images, speech, videos as well (Wang et al., 2024h)). takes the context as the input and generate a response according to the context and the model parameters. This step includes reading the extracted information from the knowledge base, the memory module, and the model parameters to provide a complete, context-aware response to the environmental query.
7.3 Summary of the Proposed Instantiation
In the above-mentioned processes, both the knowledge base and memory module are utilized, allowing us to adopt the technologies described in Saving $\mathcal{E}$ into Knowledge Bases for Retrieval and Saving $\mathcal{E}$ into Explicit Memory. Since the system incorporates a large language model, the techniques outlined in Saving $\mathcal{E}$ into Model Parameters are also applicable. During response generation, the model must handle extensive contextual inputs, which can be long enough to require the large language model to have long-context processing capabilities. By leveraging these technologies, we can enable Abstraction and Experience Merging, where the knowledge base, external memory module, and model parameters can all store abstracted versions of experiences. Then with the presence of both the knowledge base and external memory module, the system is capable of achieving Long-Term Retention and Accurate Recall via querying these modules. We would like to highlight that the current instantiation of LSCS still remains conceptual and does not represent a fully implemented system. While we do want to propose a system that can actually work, there exist some challenges, illustrated in Appendix B.
@@ 8 Conclusion
In this paper, we propose the concept of a LifeSpan Cognitive System (LSCS), which aims to manage rapid, incremental learning while retaining the ability to recall previous experiences. To achieve LSCS, we argue there are two major challenges: (1) Abstraction and Experiences Merging, (2) Long-Term Retention and Accurate Recalling. Existing technologies have the potential to partially solve these challenges. Based on our defined metric, Storage Complexity, for saving past experiences ($\mathcal{E}$), we categorize existing technologies into four distinct classes. We summarize various technologies within each category and discuss the strengths and weaknesses of the methods in each category. We argue that achieving an LSCS requires a nuanced approach integrating multiple strategies to address the above two challenges. To this end, we propose a conceptual instantiation for LSCS. We hope to provide insights into LSCS and encourage further efforts.
Acknowledgement
We thank the reviewers and the action editor for their valuable suggestions and comments. This research is based upon work supported by U.S. DARPA ECOLE Program No. HR00112390060. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Additional Details of Method Classification
Here we mainly discuss why we state that saving $\mathcal{E}$ into knowledge base has $O(n)$ storage complexity and saving $\mathcal{E}$ into memories has $o(n)$ complexity. We define Storage Complexity as a conceptual measure of how the amount of stored information grows with the number of past experiences (denoted by $n$). For example, suppose the system interacts daily with a user for one year, accumulating  365 days of conversation ($n=365$). In a knowledge-based method that stores each day’s data separately, the storage may grow linearly with the number of experiences, $O(n)$. If each day’s interaction is stored verbatim, the total storage after 365 days is approximately 365 conversation records. In contrast, memory-based methods might employ a forgetting mechanism. For instance, if the system only stores a summarized version of past interactions, it may reduce each day’s information by a certain factor. As $n$ grows large, the total amount of stored data could become sub-linear, reflecting that older interactions are increasingly compressed or partially discarded. For example, after the first few weeks, only a summary of each earlier conversation might remain. Thus, over a long period, you might not need to store all 365 full conversations; instead, you store a few full recent ones and short summaries of older ones, resulting in less than a strictly linear increase in total storage.
Appendix B Challenges of Implementing our Proposed Instantiation
Two major challenges are described below:
Limited Real-World Implementation of Memory-Based Methods: Existing memory-oriented methods for GPT-4 level large language models (LLMs) largely are essentially texts rather than true “hidden-space” memory. Methods that have memory in hidden space, such as MemoryLLM (Wang et al., 2024f), InfiniteAttention (Munkhdalai et al., 2024), Memory3 (Yang et al., 2024) illustrate early attempts in this direction. However, these methods have only been successfully applied to relatively small models (up to a few billion parameters) (Wang et al., 2024f) and often lack open-source implementations (Munkhdalai et al., 2024; Yang et al., 2024), making it difficult to scale and integrate with state-of-the-art large LLMs. In contrast, RAG methods can easily be applied on GPT4-level models. We can (1) Create an LSCS based on small models so that memory and RAG can both be incorporated, but these models may have limited capacities (For instance, RULER [4] shows that Llama-3.1-8B only has 32k effective context window.), which might make it hard to create a truly practical LSCS. (2) Ideally if we have a large memory language model where we can introduce RAG, then the built LSCS can be much stronger but as we said the "large memory language model" currently does not exist.
[…]
Despite these limitations, we view LSCS as a forward-looking framework. As larger memory-equipped models, better open-source implementations, and more extensive benchmarks emerge, the LSCS concept could transition from speculation to practical realization.
Although integrating different categories of methods might be challenging, we would like to mention that a promising direction is to use a model that supports both memory tokens and RAG in a unified manner. For instance, MemoryLLM (Wang et al., 2024f) incorporates up to 12,800 memory tokens per layer and a generation context window of 2,048 tokens. Scaling this approach—e.g., to 96k memory tokens per layer and 32k context window—could enable a model to process vast amounts of stored knowledge within its hidden space. Under such a scenario, RAG techniques could then retrieve and feed external knowledge (e.g., from a notepad-like structure) back into the model, creating a seamless interplay between internally stored representations and external references. While fine-tuning such a system remains a challenge, our view is that this would be an infrequent event. In rare cases where fine-tuning is needed, one could preserve the original training and instruction datasets, mixing them into the new training set to regularize and prevent catastrophic forgetting.